OpenStudy (anonymous):

During the same journey, Stuart drove x miles for 2 hours, and 200 miles for 3 hours. Find x if the average speed for the entire journey is 70 miles per hour

OpenStudy (anonymous):

hi

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thier result is 150 mile

OpenStudy (anonymous):

their

OpenStudy (callisto):

sorry it should be x+200 = 70*(2+3) x = 350-200 = 150

OpenStudy (callisto):

i overlooked the question.. Sorry!!!!

OpenStudy (anonymous):

it is ok

OpenStudy (anonymous):

thanks for oyur help

OpenStudy (callisto):

Distance = time x speed therefore total distance is (2+3) times 70

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

but why you use " x + 200" ?

OpenStudy (anonymous):

I don't really understand

OpenStudy (callisto):

because it is teh total distance travelled

OpenStudy (callisto):

*the

OpenStudy (anonymous):

ok I understand now.

OpenStudy (callisto):

that's good ! :)

OpenStudy (anonymous):

thanks

OpenStudy (callisto):

welcome :)

OpenStudy (anonymous):

you still there?

OpenStudy (anonymous):

Given the equations of the lines (I) 2y + 3x = 3 (II) -3y - 2x = 5 (III) -6y + 4x = 9, (IV) 2y + 6x = 9 which two lines are perpendicular? A) (I) and (II) B) (II) and (III) C) (III) and (IV) D) (I) and (III) E) (IV) and (II)

OpenStudy (anonymous):

please help me

OpenStudy (callisto):

Slope (I) -2/3 (II)-3/2 (III) 3/2 (IV)-1/3 so (I) and (III) are perpendicular answer is D good answer plz XD

OpenStudy (anonymous):

how you know (I) = -2/3?

OpenStudy (callisto):

general form of linear equation = ax+by+c =0 slope = -a/b x-int - -c/a y-int - -c/b

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

If f(x) = (x + 1)2, then f(t + 2) =?

OpenStudy (anonymous):

need help

OpenStudy (anonymous):

If you reward your helper with good answer, he/she may come back and help you again.